# Critical exponents of the Ising model with quenched structural disorder and long-range interactions at spatial dimension $d=3$

### v2

@@ Appendix B Two-loop resummation stability and three-loop resummation limitations
Let us restrict the system of equations (3.4)-(3.6) to the two-loop perturbative order and for numerical evaluation of the two-loop correlation length critical exponent $\nu$ we use the resummation techniques described in Section 4. As can be seen in Fig. 5, reliable results can be obtained even with the simplest Padé approach (blue dot-dashed line). However, our analysis shows that the Padé–Borel resummation breaks down in a certain range of the control parameter $\sigma$ (blue solid line). In contrast, the Padé–Borel–Leroy procedure remains stable (red solid line). Using the Principle of Minimal Sensitivity, we determine the optimal value $b_{opt}=27$, at which the fixed point values become stationary with respect to variations of $\sigma$.
Figure 5: Correlation length exponent $\nu$ for the Ising model at $d=3$ as a function of $\sigma$ in two-loop approximation. Dark purple dashed line: critical exponent before resummation; blue dot-dashed line: resummation using the Padé approximant; blue solid line: the Padé–Borel resummation; while red solid line: the Padé–Borel–Leroy resummation.
In comparison with the two-loop results obtained using the same resummation techniques, we are effectively unable to extract reliable estimates of the three-loop critical exponents [see Fig. 6 (a)]. While Padé approximation still shows intervals of regular behaviour separated by the region of controlled parameter where the resummed expression diverges, Padé–Borel and Padé–Borel-Leroy fail more severely, producing manifestly unphysical results. This indicates that, at the three-loop level, more sophisticated resummation strategies are required to obtain meaningful critical exponents.
A common strategy for handling asymptotically divergent series is to determine the fixed point from a resummed RG $\beta$-function at a lower loop order, and then substitute this value into the expressions for the critical exponents calculated at a higher order [65, 48, 53]. This approach typically yields more stable estimates, since $\beta$-function series are usually less well-behaved, whereas the accuracy of the critical exponents tends to improve when higher-order expansions are used [65, 48, 53]. The estimate of the correlation-length critical exponent $\nu$ obtained within this approach is shown in Fig. 6 (b).
Figure 6: Correlation length exponent $\nu$ for the Ising model at $d=3$ as a function of $\sigma$ in three-loop approximation calculated in FPs from: (a) three-loop RG $\beta$-functions and (b) two-loop RG $\beta$-functions, correspondingly. Dark purple dashed line: critical exponent before resummation; blue dot-dashed line: resummation using the Padé approximant; blue solid line: the Padé–Borel resummation; while dark purple solid line: the Padé–Borel–Leroy resummation.
@@ Appendix C Subsequent resummation technique
In our case, the RG functions contain two coupling constants: the “pure” coupling $u$ and the “disorder-induced” coupling $v$. Consequently, the RG flow is governed by two functions, $\beta_{u}(u,v)$ and $\beta_{v}(u,v)$, each of which can be expanded as a double series in powers of $u$ and $v$ [see eqs. (3.4)-(3.6)]. To improve the resummation of multivariable series, a two-step (respectively, subsequent) resummation technique based on the “inequality of rights” between the pure and random couplings was proposed in the context of the zero-dimensional ($d=0$) diluted SR Ising model [61]. This technique was then developed in the context of the three-dimensional counterpart [62, 63]. In this approach, the series are treated asymmetrically with respect to $u$ and $v$.
@@
Each coefficient $A_{k}(u)$ is itself a divergent series in $u$. For each of them, we construct the Padé-Borel-Leroy image and extrapolate it using a Padé-approximant $\left[\frac{p-k-r_{k}}{r_{k}}\right](ut)$ in the inverse transform:
$A_{k}^{(\mathrm{res})}(u)=\int_{0}^{\infty}e^{-t}\,t^{b_{1}}\,\left[\frac{p-k-r_{k}}{r_{k}}\right](ut)\,dt,$ (C.2)
where $b_{1}$ is the Leroy parameter for the $u$-series. At the calculations, we use $r_{k}=1$ for all $0\leq k\leq p$. After this stage, the function $f(u,v)$ is rewritten as
@@
$f^{(\mathrm{res})}(u,v)=\int_{0}^{\infty}e^{-s}\,s^{b_{2}}\,\left[\frac{p-1}{1}\right]_{u}(vs)\,ds,$ (C.4)
where $\left[\frac{p-1}{1}\right]_{u}(vs)$ is Padé-approximant for (C.3) with denominator lenear in $v$ and $b_{2}$ is the corresponding Leroy parameter for the $v$-series.
In practical calculations, the new parameter $q$ can be introduced to represent number of terms in (C.3) which are to be resummed. Two main variants of the subsequent resummation has been tested:
@@
    $f^{(\mathrm{res})}(u,v)=\int_{0}^{\infty}e^{-s}\,s^{b_{2}}\,\left[\frac{p-2}{1}\right]_{u}(vs)\,ds+A_{p}(u)\,v^{p}.$ (C.6)

### v3

@@ Appendix B Two-loop resummation stability and three-loop resummation limitations
Let us restrict the system of equations (3.4)–(3.6) to the two-loop perturbative order and for numerical evaluation of the two-loop correlation length critical exponent $\nu$ we use the resummation techniques described in section 4. As can be seen in figure B.1, reliable results can be obtained even with the simplest Padé approach (blue dot-dashed line). However, our analysis shows that the Padé–Borel resummation breaks down in a certain range of the control parameter $\sigma$ (blue solid line). By contrast, the Padé–Borel–Leroy procedure remains stable (red solid line). Using the Principle of Minimal Sensitivity, we determine the optimal value $b_{\rm opt}=27$, at which the fixed point values become stationary with respect to variations of $\sigma$.
Figure B.1: (Colour online) Correlation length exponent $\nu$ for the Ising model at $d=3$ as a function of $\sigma$ in two-loop approximation. Dark purple dashed line: critical exponent before resummation; blue dot-dashed line: resummation using the Padé approximant; blue solid line: the Padé–Borel resummation; while red solid line: the Padé–Borel–Leroy resummation.
In comparison with the two-loop results obtained using the same resummation techniques, we are unable to effectively extract reliable estimates of the three-loop critical exponents [see figure B.2 (a)]. While Padé approximation still shows intervals of regular behaviour separated by the region of controlled parameter where the resummed expression diverges, Padé–Borel and Padé–Borel–Leroy fail more severely, producing manifestly unphysical results. This indicates that, at the three-loop level, more sophisticated resummation strategies are required to obtain meaningful critical exponents.
A common strategy for handling asymptotically divergent series is to determine the fixed point from a resummed RG $\beta$-function at a lower loop order, and then substitute this value into the expressions for the critical exponents calculated at a higher order [65, 48, 53]. This approach typically yields more stable estimates, since $\beta$-function series are usually less well-behaved, whereas the accuracy of the critical exponents tends to improve when higher-order expansions are used [65, 48, 53]. The estimate of the correlation-length critical exponent $\nu$ obtained within this approach is shown in figure B.2 (b).
Figure B.2: (Colour online) Correlation length exponent $\nu$ for the Ising model at $d=3$ as a function of $\sigma$ in three-loop approximation calculated in FPs from: (a) three-loop RG $\beta$-functions and (b) two-loop RG $\beta$-functions, correspondingly. Dark purple dashed line: critical exponent before resummation; blue dot-dashed line: resummation using the Padé approximant; blue solid line: the Padé–Borel resummation; while dark purple solid line: the Padé–Borel–Leroy resummation.
@@ Appendix C Subsequent resummation technique
In our case, the RG functions contain two coupling constants: the “pure” coupling $u$ and the “disorder-induced” coupling $v$. Consequently, the RG flow is governed by two functions, $\beta_{u}(u,v)$ and $\beta_{v}(u,v)$, each of which can be expanded as a double series in powers of $u$ and $v$ [see equations (3.4)–(3.6)]. To improve the resummation of multivariable series, a two-step (respectively, subsequent) resummation technique based on the “inequality of rights” between the pure and random couplings was proposed in the context of the zero-dimensional ($d=0$) diluted SR Ising model [61]. This technique was then developed in the context of the three-dimensional counterpart [62, 63]. In this approach, the series are treated asymmetrically with respect to $u$ and $v$.
@@
Each coefficient $A_{k}(u)$ is itself a divergent series in $u$. For each of them, we construct the Padé–Borel–Leroy image and extrapolate it using a Padé-approximant $\big[\frac{p-k-r_{k}}{r_{k}}\big](ut)$ in the inverse transform:
$A_{k}^{(\mathrm{res})}(u)=\int_{0}^{\infty}\mathrm{e}^{-t}\,t^{b_{1}}\,\left[\frac{p-k-r_{k}}{r_{k}}\right](ut)\,\mathrm{d}t,$ (C.2)
where $b_{1}$ is the Leroy parameter for the $u$-series. At the calculations, we use $r_{k}=1$ for all $0\leqslant k\leqslant p$. After this stage, the function $f(u,v)$ is rewritten as
@@
$f^{(\mathrm{res})}(u,v)=\int_{0}^{\infty}\mathrm{e}^{-s}\,s^{b_{2}}\,\left[\frac{p-1}{1}\right]_{u}(vs)\,\mathrm{d}s,$ (C.4)
where $\big[\frac{p-1}{1}\big]_{u}(vs)$ is Padé-approximant for (C.3) with denominator linear in $v$, and $b_{2}$ is the corresponding Leroy parameter for the $v$-series.
In practical calculations, the new parameter $q$ can be introduced to represent the number of terms in (C.3) which are to be resummed. Two main variants of the subsequent resummation have been tested:
@@
    $f^{(\mathrm{res})}(u,v)=\int_{0}^{\infty}\mathrm{e}^{-s}\,s^{b_{2}}\,\left[\frac{p-2}{1}\right]_{u}(vs)\,\mathrm{d}s+A_{p}(u)\,v^{p}.$ (C.6)